Jump to main content.


Drake Chemical

EPA Daily Situation Reports for June 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Tuesday, June 30, 1998

FOR DATES: From 7:00 a.m. Monday, June 29, 1998 to 7:00 a.m. Tuesday, June 30, 1998.

WEATHER: Sunny with haze. The temperature range was 69oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 16.3 hours and 683.6 tons of contaminated soil was treated during the time frame of this report. Site soil was processed at approximately 41.9 tons per hour. Feed was halted at 1037 hours and resumed at 1819 hours for the reason noted below.

An employee working in the debris separation building suffered from heat exhaustion on Monday morning. The employee was transported to the Lock Haven Hospital, treated, and released the same day. As a result of the incident, OHM suspended operations to evaluate their heat stress monitoring program. Additional measures have been implemented to monitor workers on site. The heat stress prevention and monitoring program will continue to be evaluated by OHM and the U.S. Army Corps of Engineers.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, excavation, feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-06-23-98-147) and fly ash analytical report (fly-ash-06-23-98-025). The bottom ash met project specifications for backfill at or above elevation 547, and the fly ash met project specifications for backfill at or above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE).


Monday, June 29, 1998

FOR DATES: From 7:00 a.m. Sunday, June 28, 1998 to 7:00 a.m. Monday, June 29, 1998.

WEATHER: Sunny. The temperature range was 66oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.4 hours and 932.4 tons of contaminated soil was treated during the time frame of this report. Site soil is being processed at approximately 41.6 tons per hour. Off-line time was attributed to weight belt calibration, change out of a feed screw motor coupler, and combustion efficiency interlock caused by trippage of CEM bank B chiller. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. The contractor may resume drum-bulking activities during the night shift to take advantage of the cooler temperatures.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: D. Zion (Montgomery Watson) for David Modricker (USACE).


Sunday, June 28, 1998

FOR DATES: From 7:00 a.m. Saturday, June 27, 1998 to 7:00 a.m. Sunday, June 28, 1998.

WEATHER: Sunny. The temperature range was 65oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 50 min. and 1014.4 tons of contaminated soil was treated during the time frame of this report. . Site soil is being processed at approximately 43 tons per hour. Off-line time was attributed to weight belt calibration. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Concover was placed on the feed stockpile. Received analytical reports for bot-ash-06-19-98-143, 06-20-98-144, and 06-22-98-146. The bottom ash samples met the project specifications for backfilling at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. The contractor may resume drum-bulking activities during the night shift to take advantage of the cooler temperatures.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: R. Lupold and David Modricker(USACE)


Saturday, June 27, 1998

FOR DATES: From 7:00 a.m. Friday, June 26, 1998 to 7:00 a.m. Saturday, June 27, 1998.

WEATHER: Sunny. Humid with afternoon thunderstorms. The temperature range was 71*F to 84*F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, 7 days a week operating the Thermal Destruction Facility (TDF). The TDF was brought back on-line at 1015 hours on Friday morning. The TDF was on-line for 18.6 hours and 761 tons of contaminated soil was treated during the time frame of this report. Site soil is being processed at approximately 40.9 tons per hour. Off-line time was attributed to stopping feed during local lightning storm activity, repair of a broken feed screw drive chain, and jams in the feed screws. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling and rock washing and inorganic debris decontamination. Continued off-site and perimeter air monitoring activities. Drum bulking activities were postponed due to the heat.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: R. Lupold and David Modricker(USACE)


Friday, June 26, 1998

FOR DATES: From 7:00 a.m. Thursday, June 25, 1998 to 7:00 a.m. Friday, June 26, 1998.

WEATHER: Sunny with Haze. The temperature range was 69oF to 89oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently in a heat-up mode in preparation for resuming hazardous waste feed. The TDF has been off-line to allow for the replacement of the kiln feed screws and the feed screw jacket. The contractor continues to work 24 hours a day, seven days a week performing maintenance to the TDF equipment. At 1003 hours, a Thermal Relief Valve (TRV) opening occurred as a result of a tripped circuit breaker in the control room. Feed had been shut down for approximately 40 hours prior to the opening. The contractor notified the proper authorities and the TRV was closed at approximately 1009 hours.

II. ACTIONS TAKEN: Began bulking compatible drum contents for shipment to a hazardous waste landfill. Continued feed preparation, debris separation, and rock washing & inorganic debris decontamination. Continued the placement of Concover on the stockpile. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: The TDF is expected to resume feed on Friday morning, June 26, 1998. Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Continue drum bulking for shipment off-site.

IV. PROBLEMS NOTED OR ANTICIPATED: NONE

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Thursday, June 25, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 24, 1998 to 7:00 a.m. Thursday, June 25, 1998.

WEATHER: Sunny with Haze. The temperature range was 66oF to 84oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line to allow for replacement of the kiln feed screws and the water cooling jacket. The contractor continues to work 24 hours a day, seven days a week performing maintenance to the TDF equipment. The contractor is performing related activities such as soil feed preparation and debris separation. Operations begin today to bulk overpacked drums that are in temporary storage on-site. The drums will be combined according to chemical compatibility and shipped off-site to a hazardous waste landfill.

II. ACTIONS TAKEN: Continued excavation, feed preparation, debris separation, and rock washing & inorganic debris decontamination. Continued the placement of Concover on the stockpile. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: The TDF is expected to resume operations on Friday, June 26 or Saturday, June 27, 1998. Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Wednesday, June 24, 1998

FOR DATES: From 7:00 a.m. Tuesday, June 23, 1998 to 7:00 a.m. Wednesday, June 24, 1998.

WEATHER: Humid with afternoon rain. The temperature range was 68oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 9.0 hours and 213.3 tons of soil were treated during the time frame of this report. Off-line time was attributed to multiple feed screw stoppages and a planned shutdown to perform TDF maintenance. Site soil was processed at approximately 23.7 tons per hour.

The TDF was placed in hot standby at 1616 hours in preparation for replacement of the feed screws and cooling jacket. Maintenance performed during this period included cleaning ducts, the cyclone and the bottom of the secondary combustion chamber. A planned TRV opening occurred at1744 hours as part of the TDF system interlock checks.

II. ACTIONS TAKEN: Continued excavation, feed preparation, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: The TDF is expected to resume operations on Friday, June 26, 1998. Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Denis duBreuil, Jared Olsen (USACE), Mike Welch, Larry Newcomer, Steve Webster (PADEP).

PREPARED BY: Steve Gibson and David Modricker (USACE)


Tuesday, June 23, 1998

FOR DATES: From 7:00 a.m. Monday, June 22, 1998 to 7:00 a.m. Tuesday, June 23, 1998.

WEATHER: Partly sunny with afternoon rain. The temperature range was 68oF to 79o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours, and 620.1 tons of soil were treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and multiple feed screw stoppages. Site soil is being processed at approximately 26.3 tons per hour. The reduced feed rate is due to wear on the feed screws and material handling issues. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued excavation, feed preparation, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical reports for bot-ash-06-18-98-142 and bot-ash-06-17-98-141. Bottom ash tested met project specifications for backfill at or above elevation 547.00. Received fly ash analytical report for fly-ash-06-18-98-024. Fly ash tested met project specifications for backfill at or above elevation 553.

III. FUTURE PLANS: A shut down of the TDF is planned for Wednesday, June 24, 1998. Feed will be shut down for approximately 3 days to replace the dual feed screws and the feed screw cooling jacket. The contractor is planning on performing routine maintenance and inspections on the TDF during the shut down. Major activities for the next week also include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Monday, June 22, 1998

FOR DATES: From 7:00 a.m. Sunday, June 21, 1998 to 7:00 a.m. Monday, June 22, 1998.

WEATHER: Partly sunny. The temperature range was 65oF to 85o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours, 36 min. and 710 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and feed screw stoppages. Site soil is being processed at approximately 30 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: NONE

PREPARED BY: Danny Zion (Montgomery Watson) for David Modricker (USACE)


Sunday, June 21, 1998

FOR DATES: From 7:00 a.m. Saturday, June 20, 1998 to 7:00 a.m. Sunday, June 21, 1998.

WEATHER: Partly sunny, humid, thunderstorm. The temperature range was 68oF to 85o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22 hours, 05 min. and 711.0 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration, feed screw stoppage due to a build up of feed materials within the hopper and stoppage during a local thunderstorm. Site soil is being processed at approximately 32.2 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received ash analytical reports for bot-ash-06-16-98-140. The bottom ash test results indicated that ash met project specifications for backfill at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: NONE

PREPARED BY: Richard Lupold (USACE)


Saturday, June 20, 1998

FOR DATES: From 7:00 a.m. Friday, June 19, 1998 to 7:00 a.m. Saturday, June 20, 1998.

WEATHER: Partly cloudy to Sunny. The temperature range was 59oF to 77o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 802.1 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and to feed screw stoppage due to a build up of feed materials within the hopper. Site soil is being processed at approximately 33.4 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued excavation, drum handling, feed preparation, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received ash analytical reports for bot-ash-06-15-98-139 and fly-ash-06-15-98-023. The bottom ash test results indicated that ash met project specifications for backfill at or above elevation 547, and the fly ash test results indicated that fly ash met project specifications for backfill at or above elevation 553. AirRecon left the site on this date.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs & Mike Welch (PADEP)

PREPARED BY: Steve Gibson (USACE) and Danny Zion (Montgomery Watson).


Friday, June 19, 1998

FOR DATES: From 7:00 a.m. Thursday, June 18, 1998 to 7:00 a.m. Friday, June 19, 1998.

WEATHER: Partly cloudy to Sunny. The temperature range was 59oF to 81o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 825.7 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and to remove a buildup of feed soils within the feed screw hopper. Site soil is being processed at approximately 34.5 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. AirRecon finished stack testing Thursday evening. Continued excavation, drum handling, feed preparation, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP), Chris McBride (FOCUS Environmental), Kurt Davis (GAP).

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Thursday, June 18, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 17, 1998 to 7:00 a.m. Thursday, June 18, 1998.

WEATHER: Partly cloudy. The temperature range was 59oF to 79o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 885.7 tons of soil were treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and feed screw stoppages. Site soil is being processed at approximately 37.4 tons per hour. Feed rates varied due to wet feed material binding in the feed screws.

II. ACTIONS TAKEN: AirRecon conducted 3 runs of stack testing for dioxin and furans. Continued TDF operations. Continued feed preparation, waste soil excavation, drum handling, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical results for bot-ash-06-12-98-136 and bot-ash-06-13-98-137. Both bottom ash samples met project specifications for backfilling above elevation 547.

III. FUTURE PLANS: AirRecon will continue with stack sampling on Thursday, June 18, 1998. Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Kurt Davis (GAP), Michael Welch (PADEP)

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Wednesday, June 17, 1998

FOR DATES: From 7:00 a.m. Tuesday, June 16, 1998 to 7:00 a.m. Wednesday, June 17, 1998.

WEATHER: Overcast with Thunderstorms. The temperature range was 61oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.2 hours and 967.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to weight belt calibration, feed screw stoppages, and as a precautionary measure associated with local lightning storm events. Site soil is being processed at approximately 41.7 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical report for bot-ash-06-11-98-135. Ash met project specifications for backfill at or above elevation 547. AIR-RECON performed two runs of stack tests for metals and particulates. An additional run of stack testing for metals and particulates is planned since feed samples were not collected per the sampling plan.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs, Larry Newcomer, Michael Welch (PADEP), Chris McBride (Focus Environmental), Bob Carnein and students (Lock Haven University), and Kurt Davis (GAP).

PREPARED BY: Steve Gibson and David Modricker (USACE).


Tuesday, June 16, 1998

FOR DATES: From 7:00 a.m. Monday, June 15, 1998 to 7:00 a.m. Tuesday, June 16, 1998.

WEATHER: Overcast with periods of heavy rain. The temperature range was 56oF to 68oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 1011.7 tons of soil was treated during the time frame of this report. Off-line time was attributed to weight belt calibration. Site soil is being processed at approximately 42.3 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities. Received bottom ash analytical report for bot-ash-06-10-98-134. Ash met project specifications for backfill at or above elevation 547. AIR-RECON arrived onsite to perform stack sampling for dioxin and furans, metals and particulates.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Chris McBride (Focus Environmental)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Monday, June 15, 1998

FOR DATES: From 7:00 a.m. Sunday, June 14, 1998 to 7:00 a.m. Monday, June 15, 1998.

WEATHER: Mostly cloudy, thunderstorms. The temperature range was 56oF to 74oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 16 hours 20 minutes and 669.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to feed stoppage to repair a weigh belt splice. Site soil is being processed at approximately 42 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Danny Zion (Montgomery Watson) for David Modricker (USACE).


Sunday, June 14, 1998

FOR DATES: From 7:00 a.m. Saturday, June 13, 1998 to 7:00 a.m. Sunday, June 14, 1998.

WEATHER: Overcast with periods of heavy rain and thunderstorms. The temperature range was 58.2oF to 73.7oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 18 minutes and 979.7 tons of soil was treated during the time frame of this report. Off-line time was attributed to feed stoppage during a local lightning storm and weigh belt calibration. Site soil is being processed at approximately 42 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, WWTP operations, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring programs.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold and Steve Gibson (USACE).


Saturday, June 13, 1998

FOR DATES: From 7:00 a.m. Friday, June 12, 1998 to 7:00 a.m. Saturday, June 13, 1998.

WEATHER: Overcast with periods of heavy rain and thunderstorms. The temperature range was 64oF to 73oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 21.1 hours and 868.2 tons of soil was treated during the time frame of this report. Off-line time was attributed to feed stoppage during a local lightning storm, weigh belt calibration, brief power loss to the power screens in the feed preparation building. Site soil is being processed at approximately 41.1 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Stockpiled washed rock and concrete rubble in the lower backfill area. Continued off-site and perimeter air monitoring activities. The SODAR meteorological station was observed to be back on-line at 1000 hours.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Friday, June 12, 1998

FOR DATES: From 7:00 a.m. Thursday, June 11, 1998 to 7:00 a.m. Friday, June 12, 1998.

WEATHER: Overcast, rain Friday morning. The temperature range was approximately 45oF to 65oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.3 hours and 984.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and to repair a tear in the feed belt. Feed was stopped Friday morning due to a very short power interruption to burners and compressors. Power to the ID fan and Thermal Relief Vent was not interrupted. Incinerator operators quickly restored compressors and burners. The Thermal Relief Vent remained shut. Site soil is being processed at approximately 42.3 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Stockpiled washed rock and concrete rubble in the lower backfill area. Continued off-site and perimeter air monitoring activities. Received analytical reports for bottom ash bot-ash-06-06-98-131A and bot-ash-06-07-98-132, which met project specifications for backfill above elevation 547. Received analytical reports for fly ash, fly-ash-06-06-98-020A and fly-ash-06-06-98-21A, which met project specifications for backfill above elevation 553.

An investigation of the local SODAR station by USACE representatives revealed that the phone connection updating the SODAR information was not active at the pole. USACE contacted Weston and it is understood that Bell Atlantic maintenance crews are currently working to restore the phone connection.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Kurt Davis (GAP).

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Thursday, June 11, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 10, 1998 to 7:00 a.m. Thursday, June 11, 1998.

WEATHER: Overcast with light rain. The temperature range was 57oF to 63o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week, operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.4 hours and 931.8 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and an area grid power failure. Site soil is being processed at approximately 41.6 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, WWTP operations, debris separation, rock washing & inorganic debris decontamination, ash handling, and ash backfilling. Stockpiled washed rock and concrete rubble in the lower backfill area. Continued off-site and perimeter air monitoring activities. Excavation into the existing stockpile was performed to prepare a staging area for future excavation activities.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: A local grid power failure at approximately 0806 hours caused an unplanned Thermal Relief Valve (TRV) opening at approximately 0806, which lasted about 1 minute. Feed to the incinerator was immediately shut down and the contractor promptly brought the situation under control by notifying the proper authorities and closing the TRV. Feed resumed at approximately 0945 hours. The USACE was informed by PP&L personnel that the momentary power loss event affected only a small part of Lock Haven, and was probably caused by a PP&L maintenance crew or a biological interruption.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Wednesday, June 10, 1998

FOR DATES: From 7:00 a.m. Tuesday June 9, 1998 to 7:00 a.m. Wednesday, June 10, 1998.

WEATHER: Partly cloudy. The temperature range was approximately 45oF to 65o F.

I. CURRENT SITUATION: Heat up resumed to the Thermal Destruction Facility (TDF) in the morning and the TDF was brought back on-line Tuesday afternoon. Feed resumed at 1432 hours. The contractor is working 24 hours a day, seven days a week operating the TDF. The TDF was on-line for 16.5 hours and 679.7 tons of soil was treated during the time frame of this report. Off-line time was attributed to the repair/replacement of the lower double dump valve. Site soil is being processed at approximately 41 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, rock washing, inorganic debris decontamination, ash handling, and ash backfilling. Concover was placed on the feed stockpile for dust control and to suppress site odors. Continued off-site and perimeter air monitoring programs. Received analytical report for bot-ash-06-05-98-130. Bottom ash met project specifications for backfilling at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation and drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Warren Hauseman, Health & Safety Director (IT), Kurt Davis (GAP), Larry Newcomer, Mike Welch, and Steve Webster (PADEP).

PREPARED BY: Steve Gibson and David Modricker, (USACE).


Tuesday, June 9, 1998

FOR DATES: From 7:00 a.m. Monday, June 7, 1998 to 7:00 a.m. Tuesday, June 8, 1998.

WEATHER: Windy and partly sunny. The temperature range was approximately 41o F to 67o F.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) remains in hot standby while repairs are being made to the flap on the lower dump valve located between the dry and wet ash drag conveyors. The contractor is working 24 hours a day repairing and maintaining the TDF. It is expected that the TDF will be brought back on line Tuesday afternoon. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring. Received analytical reports for bot-ash-06-04-98-129, fly-ash-06-03-98-018A, and fly-ash-06-04-98-019. The bottom ash sample met the project specifications for backfilling at or above elevation 547 and both fly ash samples met the project specifications for backfilling at or above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF following the repair/replacement of a flap valve in the double dump valve unit, excavation and drum handling, continuation of feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination. Stack sampling for dioxins and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: During an inspection of the TDF, a water leak was observed in the feed screw cooling jacket. The contractor configured the same water-cooled system to utilize air-cooling methods. The double dump valve is being repaired as noted above.

V. NOTED VISITORS: Warren Hauseman, Health & Safety Director (IT)

PREPARED BY: Steve Gibson and David Modricker (USACE).


Monday, June 8, 1998

FOR DATES: From 7:00 a.m. Sunday, June 7, 1998 to 7:00 a.m. Monday, June 8, 1998.

WEATHER: Light rain Sunday. The temperature range was approximately 46o F to 58o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.47 hours and 849.5 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and stoppage of feed at 0632 to perform maintenance inspection of entire TDF unit. It is expected that the TDF will be off-line until Monday night to accomplish repair of the lower double dump valve and plumb the feed screw cooling jacket for air. Site soil is being processed at approximately 36 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring. Received analytical reports for bot-ash-06-02-98-127A and bot-ash-06-03-98-128. Both bottom ash samples met the project specifications for backfilling at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF following the repair/replacement of a flap valve in the double dump valve unit, excavation and drum handling, continuation of feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None noted.

PREPARED BY: Chad Thompson and David Modricker (USACE).


Sunday, June 7, 1998

FOR DATES: From 7:00 a.m. Saturday, June 6, 1998 to 7:00 a.m. Sunday, June 7, 1998.

WEATHER: Sunny to Partly Cloudy. The temperature range was approximately 43.2o F to 64.7o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours, 18 minutes and 823.7 tons of soil was treated during the time frame of this report. Off-line time was attributed to the weight belt calibration and Secondary Combustion Chamber temperature fluctuations. Site soil is being processed at approximately 35 tons per hour (tph). Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring. Received analytical reports for bot-ash-06-01-98-126. The bottom ash sample met the project specifications for backfilling at or above elevation 553. Performed maintenance on the V-320 Cyclone to remove a build up of materials with an air lance. Continued to perform external maintenance and equipment monitoring of the TDF while operating at 35 tph.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; restarting of excavation and drum handling; continuation of feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold and David Modricker (USACE).


Saturday, June 6, 1998

FOR DATES: From 7:00 a.m. Friday, June 5, 1998 to 7:00 a.m. Saturday, June 6, 1998.

WEATHER: Partly Cloudy. The temperature range was approximately 45o F to 65o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 965.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration and to remove accumulated soils on the feed screws. Site soil is being processed at approximately 40 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued off-site and perimeter air monitoring. Received analytical reports for bot-ash-05-30-98-124 and bot-ash-05-31-98-125. Both bottom ash samples met the project specifications for backfilling at or above elevation 547.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; restarting of excavation and drum handling; continuation of feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Tim Bruker & Bill Bailey (WNEP), Kurt Davis (GAP)

PREPARED BY: Steve Gibson for David Modricker (USACE).


Friday, June 5, 1998

FOR DATES: From 7:00 a.m. Thursday, June 4, 1998 to 7:00 a.m. Friday, June 5, 1998.

WEATHER: Clear. The temperature range was 42o F to 68o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 1009.6 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibration. Site soil is being processed at approximately 42.2 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Concover was placed on the feed stockpile to control dust and odors. Continued off-site and perimeter air monitoring. Received analytical reports for bot-ash-05-28-98-122, bot-ash-05-29-98-123, and fly-ash-5-29-98-017. Both bottom ash samples met the project specifications for backfilling at or above elevation 547 and the fly ash met the project specification to be backfilled at or above elevation 553.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; restarting of excavation and drum handling; continuation of feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: USACE - Roy Vaughn, Frank Meisel, Manny Diaz, Stanley Green, Brad James, Guy Jean-Baptiste

PREPARED BY: Steve Gibson for Bill Werntges, (USACE).


Thursday, June 4, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 3, 1998 to 7:00 a.m. Thursday, June 4, 1998.

WEATHER: Clear. The temperature range was 42* F to 63* F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 1011.7 tons of soil was treated during the time frame of this report. Off-line time is due to weight belt calibrations. Site soil is being processed at approximately 42.3 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued rock washing and debris decontamination. Continued off-site and perimeter air monitoring. Received bottom ash analytical for bin #7 (BOT-ASH-5-27-98-122), ash was approved for backfilling at elevation 547. A site tour was given to Lock Haven University students. A junior partnering meeting between the USACE, OHM, and PADEP was held to address junior staff concerns at the Drake site.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Mike Welch (PADEP)

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Wednesday, June 3, 1998

FOR DATES: From 7:00 a.m. Tuesday, June 2, 1998 to 7:00 a.m. Wednesday, June 3, 1998.

WEATHER: Clear. The temperature range was 45* F to 79* F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.05 hours and 927.4 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibrations and feed stoppage due to local electrical storm. Site soil is being processed at approximately 42.1 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued WWTP operations, feed preparation, debris separation, ash handling, and ash backfilling. Continued rock washing and debris contamination. Continued off-site and perimeter air monitoring.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs, Mike Welch, (PADEP); Gregg Crystall, John Banks, (USEPA)

PREPARED BY: Steve Gibson for Bill Werntges, (USACE).


Tuesday, June 2, 1998

FOR DATES: From 7:00 a.m. Monday, June 1, 1998 to 7:00 a.m. Tuesday, June 2, 1998.

WEATHER: Clear. The temperature range was 45oF to 71o F.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 1006.4 tons of soil was treated during the time frame of this report. Off-line time was attributed to weigh belt calibrations. Site soil is being processed at approximately 42.3 tons per hour. Related ongoing activities include soil feed preparation and ash handling.

II. ACTIONS TAKEN: Continued TDF operations. Continued feed preparation, debris separation, ash handling, and ash backfilling. Placing imported backfill from elevation 545 to 547 in backfill area near station 3+50. Continued off-site and perimeter air monitoring. Drum carcasses and PPE were transferred off-site to a hazardous waste landfill.

III. FUTURE PLANS: Major activities for the next week include: operation of the TDF; continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock washing & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for Frank Cleaver, (USACE).


Monday, June 1, 1998

FOR DATES: From 7:00 a.m. Sunday, May 31, 1998 to 7:00 a.m. Monday, June 1, 1998.

WEATHER: Mostly cloudy. Thunderstorm and rain in evening. The temperature range was 53oF to 86o F.

I. CURRENT

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.